2. Express the following linear equati.ons in the form a.x + by
+ c = 0 and indica.U! the values of a, b and c in each case:
(i) 2x + 3y = 9.36 (ii) x - - 10 = 0 (iii) -2x + 3y = 6
(iv) x = 3y (v) 2x = - 5y (vi) 3x + 2 = 0
(vii) y - 2 = 0 (viii) 5 = 2x.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
2x + 3y = 9. - 2x + 3y - 9. - = o. Here, a = 2, b = 3, c = - 9.35. (ii) x - f - 10 = 0, here a = 1, b = - l and c = - 10. (iii) - 2x + 3y = 6 - 2x + 3y - 6 = 0. Here, a = - 2, b = 3, c = - 6. (iv) x = 3y 1x - 3y + 0 = 0. Here, a = 1, b = - 3, c = 0. (v) 2x = - 5y 2x + 5y + 0 = 0. Here, a = 2, b = 6, c = 0. (vi) 3x + 2 = 0 3x + Oy + 2 = 0. Here, a = 3, b = 0, c = 2. (vii") y - 2 = 0 O.x + 1.y - 2 = 0. Here, a = 0, b = 1, c = - 2. (viii) 5 = 2x 2x + O.y - 5 = O. Here, a = 2, b = 0, c = - 5.